# Coupled Microelectromechanical Drum Resonators for Reservoir Computing via Sideband Pumped Phonon-Cavity Dynamics

### v1

@@ Abstract
Reservoir computing is a bio-inspired machine learning paradigm that exploits the intrinsic dynamics of nonlinear systems with fading memory for efficient temporal information processing. Microelectromechanical resonators offer a promising platform for reservoir computing as they inherently possess the requisite nonlinear and temporal properties while also facilitating the integration of sensing and computing within a single platform. In this work, we experimentally demonstrate a physical reservoir computing platform based on two capacitively coupled drum resonators, operating in the MHz frequency regime. Taking advantage of the concept of phonon-cavity electromechanics, a pump tone is applied at the sideband of the phonon cavity while probing one of the coupled modes, analogous to optomechanical systems, thereby creating nonlinear dynamics in energy transfer between the two resonators. Physical reservoir computing is implemented by exploiting the nonlinear response induced through pump amplitude modulation in combination with a time-delay feedback loop, and the performance is evaluated using both parity and Normalized Auto-Regressive Moving Average benchmarks. This work demonstrates a compact microelectromechanical platform for the integration of sensing and reservoir computing. Moreover, the sideband pumping scheme can further extend conventional single resonator reservoir computing to a multimode architecture.
@@ 1 Introduction
Reservoir Computing is a bio-inspired machine learning paradigm designed for processing complex time-series data [RC_overview, jaeger, maass]. It has been successfully applied to a diverse range of temporal tasks, including chaotic time-series prediction, speech recognition, and signal classification. As a simplified framework derived from Recurrent Neural Networks (RNNs) [graves, werbos], reservoir computing fixes the internal weights of the reservoir and only requires the output layer weights to be trained by using simple linear regression. It therefore significantly reduces computational costs while circumventing the vanishing gradient problems typical of traditional RNN training. A physical system used to implement a reservoir must exhibit two key properties: nonlinearity and fading memory [samarasinghe, appeltant]. Nonlinearity is required to project input signals into a high-dimensional state space, enabling complex data to become linearly separable at the readout layer. Fading memory is essential for processing temporal sequences, as it establishes a connection between the current reservoir state and recent past inputs. To date, reservoir computing has been demonstrated across a wide range of hardware platforms, including photonic and optoelectronic systems [optoelectronic_RC_1, optoelectronic_RC_2, optics_RC], robotics [mechanics_RC_2, mechanics_RC_1], memristor arrays [memristors_RC], spintronic [spintronic_RC, spin_waves_2, RC_spin_waves]. More recently, reservoir computing has been further extended beyond classical systems to quantum platforms, including quantum oscillators and nonlinear oscillators coupled to quantum systems [dudas2023quantum, govia2021quantum].
@@
Micro-Electro-Mechanical Systems (MEMS), which enable mechanical degrees of freedom to be coupled with electrical signals, constitute a promising platform for reservoir computing [Sylvestre, coulombe, guo2024mems, Sun2021, shima2025improved]. MEMS naturally possess the two essential properties required for reservoir computing: intrinsic nonlinearity arising from geometric and material constraints (such as the Duffing nonlinearity) effects, and fading memory facilitated by the finite decay time of their mechanical resonances [bachtold2022mesoscopic, cattiaux, maillet2018measuring]. In addition, MEMS devices can be designed as scalable architectures and are easily integrated with modern CMOS (Complementary Metal–Oxide–Semiconductor) platforms [fedder2015resonant]. Beyond these advantages, one of the most compelling motivations for developing MEMS-based reservoir computing is its potential to combine sensing and computing within a single platform [nikfarjamenergy, bai2025mems]. Such integration will overcome the inefficiencies of conventional architectures, where the physical separation of sensing and processing units leads to additional energy consumption and latency due to intensive data transfer. To date, MEMS-based physical reservoir computing has been experimentally implemented primarily using single electromechanical resonators, exploiting driving-force modulation in the Duffing nonlinear regime [Sylvestre] and stiffness modulation techniques [guo2024mems]. Only a limited number of studies have explored reservoir computing in coupled MEMS systems. These efforts include numerical investigations of reservoir performance in coupled resonators [zheng2021enhancing] and experimental studies evaluating memory capacity in coupled triple-resonator configurations [shima2025improved]. However, they rely on mechanical coupling and require the resonators to have closely matched resonance frequencies, which places stringent demands on nanofabrication and poses challenges for scaling to higher integration densities.
@@
In this work, we experimentally demonstrate a novel reservoir computing scheme based on coupled microelectromechanical resonators by exploiting concepts from phonon-cavity electromechanics [mahboob2012phonon, sun2016correlated]. The device employed is a double-drum electromechanical system consisting of two capacitively coupled membrane resonators operating in the MHz frequency regime [xin_zhou, alok]. Drawing an analogy to two-tone operation in optomechanical systems [kumar2023microwave], we apply a pump tone at the sideband of the phonon-cavity while probing one of the coupled modes, thereby inducing nonlinear dynamics in the phonon transfer between the two vibrational modes [alok]. By modulating the pump amplitude and implementing a time-delay feedback loop, we demonstrate reservoir computing by probing one of the coupled drum resonators and evaluating both parity and NARMA benchmarks. The unique device design and pump amplitude modulation scheme point toward a compact MEMS platform for co-integrated sensing and reservoir computing. Moreover, the proposed scheme does not require the resonance frequencies to fall within a narrowly distributed range and is not limited to the specific double-drum design. It can be easily extended to other mechanically coupled resonator arrays and multimode optomechanical systems, in both the classical and quantum regimes.
@@ 2.1 Double-drum resonator and setup for reservoir computing
Figure 1: Experiment setup. (a) Optical image of the Al drum resonator which is suspended over the SiN circular drum by the means of 4 support feet. (b) Optical image of the bottom SiN membrane resonator, covered with an Al thin film. There is no physical connection between the two drums. (c) Cross sectional view of the electromechanical system. (d) and (f) The mechanical responses of the SiN and Al drum resonators are measured under a $DC$ bias of $V_{dc}$ = 4 V, with ACAC drive of $V_{d}$ = 6 mV and 80 mV, respectively, each attenuated by 20 dB. (e) Schematic diagram of the measurement setup. The mechanical displacement is read out using a lock-in amplifier via microwave optomechanical interferometry. The output of the FPGA is used to modulate the pump signal through a multiplier.
@@
The microelectromechanical system used to perform reservoir computing in this work is a double-drum resonator, consisting of two suspended membrane electromechanical resonators, as shown in Fig.1(a)–(c). A suspended aluminum (Al) membrane serves as a capacitively coupled top gate to the underlying silicon nitride (SiN) drum microelectromechanical resonator [xin_zhou, alok]. In order to excite mechanical vibrations, electrostatic forces are generated by applying a combination of a $DC$ voltage $V_{dc}$ and an $AC$ voltage $V_{ac}$. We exploit the microwave optomechanical interferometry to simultaneously readout the vibrations of both drums, as shown in Fig.1(e). Microwave photons at a frequency $\omega_{c}$ = 6 GHz are shined to the double-drum resonator through a 50 Ohm transmission line. The reflected microwave signal, carrying information of the mechanical vibrations at frequency $\omega_{c}+\Omega$, is readout by using a lock-in amplifier after frequency down-conversion [xin_zhou]. Mechanical mode vibrations are therefore measured in volts in our experiment. Using this readout scheme, the double-drum resonator is initially characterized in its linear operating regime. The SiN drum has a resonance frequency of $\Omega_{SiN}/(2\pi)\approx$ 12.265 MHz and a quality factor of $Q_{SiN}\approx$ 1.3 $\times 10^{4}$. The Al drum resonates at the frequency $\Omega_{Al}/(2\pi)\approx$ 6.13 MHz with a quality factor $Q_{Al}\approx$ 237. All measurements are carried out at room temperature and under vacuum conditions, $\approx 6\times 10^{-6}$ mbar.
@@
The reservoir used for computing is constructed from virtual nodes based on a single double-drum resonator acting as the physical node. The creation of virtual nodes are inspired by the conventional time domain multiplexing method implemented in previous works [Feedbback, Sylvestre]. A standard time-delay loop is implemented using a field-programmable gate array (FPGA), which adds a time delay $\tau$ to the detected mechanical vibration $x(t)$ from the output of the nonlinear element and feeds it back to the input of the resonator in the form of driving forces. The virtual nodes are created by mapping a predefined mask $m(t)$ onto the input data $u(t)$. The masked input is then multiplexed with the delayed output of the physical element, $x(t-\tau)$, and combined through a multiplier to generate additional modulations on the driving forces, acting on the double-drum resonator.
@@ 2.2 Nonlinear behavior generated by sideband pumping in phonon-cavity electromechanics
The key concept of phonon-cavity electromechanics is to manipulate coherent energy transfer between two coupled mechanical vibration modes, $\Omega_{1}$ and $\Omega_{2}$, by selecting the mode having the higher resonance frequency ($\Omega_{1}>\Omega_{2}$) and pumping it at its sideband $\Omega_{1}\pm\Omega_{2}$ [mahboob2012phonon, sun2016correlated, alok, xu2024imaging]. It inherits rich physics from optomechanics, enabling cooling of mechanical vibration modes, amplification of phonon occupancy, and the creation of controllable interference effects [zhou2021electric]. In this double-drum resonator, we choose the SiN drum as the phonon-cavity mode and pump it at its blue sideband, at the frequency $\Omega_{p}=\Omega_{1}+\Omega_{2}+\Delta=\Omega_{Al}+\Omega_{SiN}+\Delta$. Here the $\Delta$ is defined as small frequency detuning. Additional to the pump tone, a second tone with small amplitudes is introduced to probe one of the mode with a small frequency detuning $\delta$ from the resonance, either $\Omega_{d}\approx\Omega_{1}+\delta$ or $\Omega_{d}\approx\Omega_{2}+\delta$, placing the double-drum under a two-tone driving scheme, as shown in Fig. 2 (a) and (b). Because of frequency mixing effects, the probe and pump tones generate vibration phonons in the unprobed mode, as indicated by the process <1><1>. These phonons are subsequently projected back onto the probed mode by the pump, where they interfere with the probe tone.
Figure 2: Diagram of the blue sideband pumping scheme while probing (a) the phonon-cavity SiN drum and (b) the Al membrane resonator, respectively. (c) The measured normalized response mechanical responses as a function of the pump amplitude. The background contribution has been subtracted using the measured minimum amplitude as a reference. The black (red) curve is measured by probing Al (SiN) drum at $\Omega_{d}/(2\pi)$ = 6.12 MHz ( = 12.25 MHz) with amplitude $V_{d}$ = 80 mV (6 mV) attenuated by 20 dB, with $\Omega_{p}/(2\pi)$ = 18.38 MHz and $V_{dc}$ = 4 V. (d) Analytical calculation results of the mechanical amplitude as a function of the pump amplitude in both probing cases, by using experimental parameters.
@@
where index 1 and 2 refer to each of the SiN and Al membrane respectively. The $m_{1,2}$ denotes the effective mass of the resonator, $d$ the distance between two drums, $X_{1,2}$ the mechanical displacement, $C_{g0}$ coupling capacitance between two drums at $X_{1,2}$=0, and $\gamma$ the damping rate. Within the capacitive coupling scheme, the electrostatic force acting on the resonator arises from the term $V_{dc}V_{ac}\partial C_{g}(x_{1,2})/\partial x_{1,2}$ [xin_zhou]. In the two-tone driving scheme, $V_{ac}$ is given by $V_{p}\cos(\Omega_{p}t)+V_{d}\cos(\Omega_{d}t)$. In the case of probing the SiN drum resonator, $\Omega_{d}$ = $\Omega_{1}+\delta$ = $\Omega_{SiN}+\delta$, the displacement of the probed resonator can be obtained by solving the Eq.1. To do so, we re-write variables in complex form, such as $X_{1}(t)=\frac{x_{1}(t)}{2}e^{-i\Omega_{d}t}+c.c.$, and $x_{1}$ arrives
@@
where $f_{d}=\frac{C_{g0}V_{dc}\mu_{d}}{d}$ and $f_{p}=\frac{C_{g0}V_{dc}\mu_{p}}{d}$ are the amplitudes of the probe and pump forces respectively, the parameter $\mu$ represents the complex amplitude of the pump or probe tone, and the $c.c.$ denotes the complex conjugate term. The $\chi_{1}=\frac{1}{-\delta-i\frac{\gamma_{1}}{2}}$ and $\chi_{2}=\frac{1}{\delta-\Delta-i\frac{\gamma_{2}}{2}}$ are the mechanical susceptibilities of the drum resonators when probing the SiN phonon-cavity. The entire derivation process has been reported in our previous work [alok]. The calculation results corresponding to the case of probing the Al drum are provided in the Supporting Information (SI). From the displacement expressions of both drums, $x_{1}$ and $x_{2}$, as given in Eq.2, the blue sideband pump brings nonlinear amplifications of the phonons for the probed mode through creating frequency dependent constructive interference. Such interference arises from pump induced coherent phonon cycling between the two coupled modes, and the effective constructive interferences windows depend on the frequency detunings $\delta$ and $\Delta$ relative to the mechanical damping rates $\gamma_{1}$ and $\gamma_{2}$. The pump tone functions as a phonon bus, transferring the energy from one mode to the other, in the coupled resonators.
@@
To take advantage of this pump induced nonlinear behavior for reservoir computing, we exploit the fact that the sideband pump amplitude has a effect on shifting $\Omega_{SiN}$, so called optical spring effect in optomechanics [alok, zhou2021electric]. More details can be found in SI part. This effect arises from the fact that the sideband pump amplitude modifies the effective susceptibility of the coupled mechanical resonators. It is worth emphasizing that this pump induced effect on $\Omega_{Al}$ can be neglected in this double-drum system. Because the Al drum has a relatively large bandwidth (damping rate), $\gamma_{Al}\approx 2\pi\times 25.9,\text{kHz}$, small variations in the resonance frequency are not easily observed. To do so, we set the pump frequency to $\Omega_{p}=\Omega_{SiN}+\Omega_{Al}$, where both resonances $\Omega_{SiN}$ and $\Omega_{Al}$ are obtained by measuring mechanical responses at the sideband pump amplitude $V_{pc}$ = 105 mV. We then set the probe tone frequency to either $\Omega_{d}=\Omega_{SiN}$ or $\Omega_{d}=\Omega_{Al}$ + a few kHz, and measure its amplitude as a function of $V_{p}$. The detected probe signals exhibit clear nonlinear constructive interference induced by variations in the pump amplitude, as shown in Fig.2 (c). When the pump amplitude deviates far from the calibrated value $V_{pc}$ = 105 mV, the interference between the probe tone and the phonons projected back from the unprobed mode becomes less effective due to shifts in $\Omega_{SiN}$, which induce frequency detunings $\delta$ and $\Delta$ in Eq.2. Because the two drum resonators have very different damping rates, the pump induced nonlinear behaviors are quite different when probing the cavity mode versus the Al drum. The bandwidth $\gamma_{SiN}\approx 2\pi\times 944,\text{Hz}$ is much smaller than the resonance-frequency shift induced by variations in the pump amplitude (see SI). Consequently, when the probe tone is initially biased at $\Omega_{SiN}$ corresponding to $V_{pc}=105,\text{mV}$, both increases and decreases in the pump amplitude can shift $\Omega_{SiN}$ over a frequency rang of 7 kHz, making the probe tone to move out of resonance. However, the probe tone can still be amplified through constructive interference, since this frequency shift lies within $\gamma_{Al}$. When probing the Al drum, although the probe tone always remains within the bandwidth, clear constructive interference can be observed only when the resonance shift relative to the value calibrated at $V_{pc}$ lies within $\gamma_{SiN}$. We also noticed that the detected amplitudes of Al drum exhibiting small fluctuations for $V_{p}<$ 75 mV. It may be induced by fluctuations in the background noise carried by the pump tone. Fig. 2 (d) presents the calculation results, based on Eq.2, including the pump induced shift of $\Omega_{SiN}$, which are consistent with the experimental observations. These results demonstrate that pump amplitude modulation in the two-tone driving scheme provides a controllable way of inducing nonlinear behavior in the probe tone, in coupled resonators. In the following part, we exploit this nonlinear dynamics for reservoir computing.

### v2

@@ Abstract
Reservoir computing is a bio-inspired machine learning paradigm that exploits the intrinsic dynamics of nonlinear systems with fading memory for efficient temporal information processing. Microelectromechanical resonators offer a promising platform for reservoir computing as they inherently possess the requisite nonlinear and temporal properties while also facilitating the integration of sensing and computing within a single platform. In this work, we experimentally demonstrate a physical reservoir computing platform based on two capacitively coupled drum resonators, operating in the MHz frequency regime. Taking advantage of the concept of phonon-cavity electromechanics, a pump tone is applied at the sideband of the phonon cavity while probing one of the coupled modes, analogous to optomechanical systems, thereby creating nonlinear dynamics in energy transfer between the two resonators. Reservoir computing is implemented by exploiting the nonlinear response generated through pump amplitude modulation in combination with a time-delay feedback loop, and the performance is evaluated using both parity and Normalized Auto-Regressive Moving Average benchmarks. This work demonstrates a compact microelectromechanical platform for integrated sensing and reservoir computing and shows that the sideband pumping scheme provides a pathway for extending conventional single-resonator reservoir computing toward multimode architectures.
@@ 1 Introduction
Reservoir Computing is a bio-inspired machine learning paradigm designed for processing complex time-series data [21, 18, 22]. It has been successfully applied to a diverse range of temporal tasks, including chaotic time-series prediction, speech recognition, and signal classification. As a simplified framework derived from Recurrent Neural Networks (RNNs) [16, 37], reservoir computing fixes the internal weights of the reservoir and only requires the output layer weights to be trained by using simple linear regression. It therefore significantly reduces computational costs while circumventing the vanishing gradient problems typical of traditional RNN training. A physical system used to implement a reservoir must exhibit two key properties: nonlinearity and fading memory [29, 1]. Nonlinearity is required to project input signals into a high-dimensional state space, enabling complex data to become linearly separable at the readout layer. Fading memory is essential for processing temporal sequences, as it establishes a connection between the current reservoir state and recent past inputs. To date, reservoir computing has been demonstrated across a wide range of hardware platforms, including photonic and optoelectronic systems [27, 20, 34], robotics [10, 6], memristor arrays [12], spintronic [33, 36, 25]. More recently, reservoir computing has been further extended beyond classical systems to quantum platforms, including quantum oscillators and nonlinear oscillators coupled to quantum systems [13, 15].
@@
Micro-Electro-Mechanical Systems (MEMS), which enable mechanical degrees of freedom to be coupled with electrical signals, constitute a promising platform for reservoir computing [11, 8, 17, 32, 30]. MEMS naturally possess the two essential properties required for reservoir computing: intrinsic nonlinearity arising from geometric and material constraints (such as the Duffing nonlinearity) effects, and fading memory facilitated by the finite decay time of their mechanical resonances [3, 7, 24]. In addition, MEMS devices can be designed as scalable architectures and are easily integrated with modern CMOS (Complementary Metal–Oxide–Semiconductor) platforms [14]. Beyond these advantages, one of the most compelling motivations for developing MEMS-based reservoir computing is its potential to combine sensing and computing within a single platform [26, 4]. Such integration will overcome the inefficiencies of conventional architectures, where the physical separation of sensing and processing units leads to additional energy consumption and latency due to intensive data transfer. To date, MEMS-based physical reservoir computing has been experimentally implemented primarily using single electromechanical resonators, exploiting driving-force modulation in the Duffing nonlinear regime [11] and stiffness modulation techniques [17]. Only a limited number of studies have explored reservoir computing in coupled MEMS systems. These efforts include numerical investigations of reservoir performance in coupled resonators [41] and experimental studies evaluating memory capacity in coupled triple-resonator configurations [30]. However, they rely on mechanical coupling and require the resonators to have closely matched resonance frequencies, which places stringent demands on nanofabrication and poses challenges for scaling to higher integration densities.
@@
In this work, we experimentally demonstrate a novel reservoir computing scheme based on coupled microelectromechanical resonators by exploiting concepts from phonon-cavity electromechanics [23, 31]. The device employed is a double-drum electromechanical system consisting of two capacitively coupled membrane resonators operating in the MHz frequency regime [44, 28]. Drawing an analogy to two-tone operation in optomechanical systems [19], we apply a pump tone at the sideband of the phonon-cavity while probing one of the coupled modes, thereby inducing nonlinear dynamics in the phonon transfer between the two vibrational modes [28]. By modulating the pump amplitude and implementing a time-delay feedback loop, we demonstrate reservoir computing by probing one of the coupled drum resonators and evaluating both parity and Normalized Auto-Regressive Moving Average (NARMA) benchmarks. The unique device design and pump amplitude modulation scheme point toward a compact MEMS platform for co-integrated sensing and reservoir computing. Moreover, the proposed scheme does not require the resonance frequencies to fall within a narrowly distributed range and is not limited to the specific double-drum design. It can be easily extended to other mechanically coupled resonator arrays and multimode optomechanical systems, in both the classical and quantum regimes.
@@ 2.1 Double-drum resonator and setup for reservoir computing
Figure 1: Experiment setup. (a) Optical image of the Al drum resonator which is suspended over the SiN circular drum by the means of 4 support feet. (b) Optical image of the bottom SiN membrane resonator, covered with an Al thin film. There is no physical connection between the two drums. (c) Cross sectional view of the electromechanical system. (d) and (f) The mechanical responses of the SiN and Al drum resonators are measured under a $DC$ bias of $V_{dc}$ = 4 V, with ACAC drive of $V_{d}$ = 6 mV and 80 mV, respectively, each attenuated by 20 dB. The red lines represent the Lorentz fitting result. (e) Schematic diagram of the measurement setup. The mechanical displacement is read out using a lock-in amplifier via microwave optomechanical interferometry. The output of the FPGA is used to modulate the pump signal through a multiplier.
@@
The microelectromechanical system used to perform reservoir computing in this work is a double-drum resonator, consisting of two suspended membrane electromechanical resonators, as shown in Fig.1(a)–(c). A suspended aluminum (Al) membrane serves as a capacitively coupled top gate to the underlying silicon nitride (SiN) drum microelectromechanical resonator [44, 28]. In order to excite mechanical vibrations, electrostatic forces are generated by applying a combination of a $DC$ voltage $V_{dc}$ and an $AC$ voltage $V_{ac}$ at a drive frequency $\Omega_{d}$ close to the mechanical resonance frequency $\Omega_{m}$ of the target mechanical resonator. We exploit the microwave optomechanical interferometry to simultaneously readout the vibrations of both drums, as shown in Fig.1(e). Microwave photons at a frequency $\omega_{c}$ = 6 GHz are shined to the double-drum resonator through a 50 Ohm transmission line. The reflected microwave signal, carrying information of the mechanical vibrations at frequency $\omega_{c}+\Omega_{d}$, is readout by using a lock-in amplifier after frequency down-conversion [44]. Mechanical mode vibrations are therefore measured in volts in our experiment. Using this readout scheme, the double-drum resonator is initially characterized in its linear operating regime by applying probe tones around $\Omega_{m}$ with small $AC$ amplitudes $V_{d}$, as shown in Fig.1 (d)-(f). The SiN drum has a resonance frequency of $\Omega_{SiN}/(2\pi)\approx$ 12.265 MHz and a quality factor of $Q_{SiN}\approx$ 1.3 $\times 10^{4}$. The Al drum resonates at the frequency $\Omega_{Al}/(2\pi)\approx$ 6.13 MHz with a quality factor $Q_{Al}\approx$ 237. All measurements are carried out at room temperature and under vacuum conditions, $\approx 6\times 10^{-6}$ mbar.
@@
The reservoir used for computing is constructed from virtual nodes based on a single double-drum resonator acting as the physical node. The creation of virtual nodes is inspired by the conventional time domain multiplexing method implemented in previous works [2, 11]. A standard time-delay loop is implemented using a field-programmable gate array (FPGA), which adds a time delay $\tau$ to the detected mechanical vibration $x(t)$ from the output of the nonlinear element and feeds it back to the input of the resonator in the form of driving forces. The virtual nodes are created by mapping a predefined mask $m(t)$ onto the input data $u(t)$. The masked input is then multiplexed with the delayed output of the physical element, $x(t-\tau)$, and combined through a multiplier to generate additional modulations on the driving forces, acting on the double-drum resonator.
@@ 2.2 Nonlinear behavior generated by sideband pumping in phonon-cavity electromechanics
The key concept of phonon-cavity electromechanics is to manipulate coherent energy transfer between two coupled mechanical vibration modes, $\Omega_{1}$ and $\Omega_{2}$, by selecting the mode having the higher resonance frequency ($\Omega_{1}>\Omega_{2}$) and pumping it at its sideband $\Omega_{1}\pm\Omega_{2}$ [23, 31, 28, 40]. It inherits rich physics from optomechanics, enabling cooling of mechanical vibration modes, amplification of phonon occupancy, and the creation of controllable interference effects [43]. In this double-drum resonator, we choose the SiN drum as the phonon-cavity mode and pump it at its blue sideband, at the frequency $\Omega_{p}=\Omega_{1}+\Omega_{2}+\Delta=\Omega_{Al}+\Omega_{SiN}+\Delta$. Here the $\Delta$ is defined as small frequency detuning. Additional to the pump tone, a second tone with small amplitudes is introduced to probe one of the mode with a small frequency detuning $\delta$ from the resonance, either $\Omega_{d}\approx\Omega_{1}+\delta$ or $\Omega_{d}\approx\Omega_{2}+\delta$, placing the double-drum under a two-tone driving scheme, as shown in Fig. 2 (a) and (b). Because of frequency mixing effects, the probe and pump tones generate vibration phonons in the unprobed mode, as indicated by the process <1><1>. These phonons are subsequently projected back onto the probed mode by the pump, where they interfere with the probe tone, as marked by the process <2><2>.
Figure 2: Diagram of the blue sideband pumping scheme while probing (a) the phonon-cavity SiN drum and (b) the Al membrane resonator, respectively. (c) The measured normalized mechanical responses as a function of the pump amplitude. The background contribution has been subtracted using the measured minimum amplitude as a reference. The black (red) curve is measured by probing Al (SiN) drum at $\Omega_{d}/(2\pi)$ = 6.12 MHz ( = 12.25 MHz) with amplitude $V_{d}$ = 80 mV (6 mV) attenuated by 20 dB, with $\Omega_{p}/(2\pi)$ = 18.38 MHz and $V_{dc}$ = 4 V. (d) Analytical calculation results of the mechanical amplitude as a function of the pump amplitude in both probing cases, by using experimental parameters. (e) Schematic of the concept for implementing reservoir computing in coupled drum resonators.
@@
where index 1 and 2 refer to each of the SiN and Al membrane respectively. The $m_{1,2}$ denotes the effective mass of the resonator, $d$ the distance between two drums, $X_{1,2}$ the mechanical displacement, $C_{g0}$ coupling capacitance between two drums at $X_{1,2}$ = 0, and $\gamma$ the damping rate. Within the capacitive coupling scheme, the electrostatic force acting on the resonator arises from the term $V_{dc}V_{ac}\partial C_{g}(x_{1,2})/\partial x_{1,2}$ [44]. In the two-tone driving scheme, $V_{ac}$ is given by $V_{p}\cos(\Omega_{p}t)+V_{d}\cos(\Omega_{d}t)$. In the case of probing the SiN drum resonator, $\Omega_{d}$ = $\Omega_{1}+\delta$ = $\Omega_{SiN}+\delta$, the displacement of the probed resonator can be obtained by solving the Eq.1. To do so, we re-write variables in complex form, such as $X_{1}(t)=\frac{x_{1}(t)}{2}e^{-i\Omega_{d}t}+c.c.$, and $x_{1}$ arrives
@@
by looking for the solution of $x_{1}$ around the probe frequency $\Omega_{d}$. Here, $f_{d}=\frac{C_{g0}V_{dc}\mu_{d}}{d}$ and $f_{p}=\frac{C_{g0}V_{dc}\mu_{p}}{d}$ are the amplitudes of the probe and pump forces respectively, the parameter $\mu_{d}$ and $\mu_{p}$ represent the complex amplitude of the probe and pump tone respectively, and the $c.c.$ denotes the corresponding complex conjugate term. The $\chi_{1}=\frac{1}{-\delta-i\frac{\gamma_{1}}{2}}$ and $\chi_{2}=\frac{1}{\delta-\Delta-i\frac{\gamma_{2}}{2}}$ are the mechanical susceptibilities of the drum resonators when probing the SiN phonon-cavity. The solutions of the coupled motion equations shown in the Eq.2 clearly indicate that the phonons in the unprobed mode, here the $x_{2}$, are generated by both the probe and pump tones through frequency conversion between the $f_{d}$ and the $f_{p}$. The entire derivation process has been reported in our previous work [28]. The calculation results corresponding to the case of probing the Al drum are provided in the Supporting Information (SI). From the displacement expressions of both drums, $x_{1}$ and $x_{2}$, as given in Eq.2, the blue sideband pump brings nonlinear amplifications of the phonons for the probed mode through creating frequency dependent constructive interference. Such interference arises from pump induced coherent phonon cycling between the two coupled modes, and the effective constructive interferences windows depend on the frequency detunings $\delta$ and $\Delta$ relative to the mechanical damping rates $\gamma_{1}$ and $\gamma_{2}$. The pump tone functions as a phonon bus, transferring the energy from one mode to the other, in the coupled resonators.
@@
To take advantage of this pump induced nonlinear behavior for reservoir computing, we exploit the fact that the sideband pump amplitude has an effect on shifting $\Omega_{SiN}$, so called optical spring effect in optomechanics [28, 43]. More details can be found in SI part. This effect arises from the fact that the sideband pump amplitude modifies the effective susceptibility of the coupled mechanical resonators. It is worth emphasizing that this pump induced effect on $\Omega_{Al}$ can be neglected in this double-drum system. Because the Al drum has a relatively large bandwidth (damping rate), $\gamma_{Al}\approx 2\pi\times$ 25.9 kHz, small variations in the resonance frequency are not easily observed. To do so, we set the pump frequency to $\Omega_{p}=\Omega_{SiN}+\Omega_{Al}$, where both resonances $\Omega_{SiN}$ and $\Omega_{Al}$ are obtained by measuring mechanical responses at the sideband pump amplitude $V_{pc}$ = 105 mV. We then set the probe tone frequency to either $\Omega_{d}=\Omega_{SiN}$ or $\Omega_{d}=\Omega_{Al}$ + a few kHz, and measure its amplitude as a function of $V_{p}$. The detected probe signals exhibit clear nonlinear constructive interference induced by variations in the pump amplitude, as shown in Fig.2 (c). When the pump amplitude deviates far from the calibrated value $V_{pc}$ = 105 mV, the interference between the probe tone and the phonons projected back from the unprobed mode becomes less effective due to shifts in $\Omega_{SiN}$, which induce frequency detunings $\delta$ and $\Delta$ in Eq.2. Because the two drum resonators have very different damping rates, the pump induced nonlinear behaviors are quite different when probing the cavity mode versus the Al drum. The bandwidth $\gamma_{SiN}\approx 2\pi\times$ 944 Hz is much smaller than the resonance-frequency shift induced by variations in the pump amplitude (see SI). Consequently, when the probe tone is initially biased at $\Omega_{SiN}$ corresponding to $V_{pc}$ = 105 mV, both increases and decreases in the pump amplitude can shift $\Omega_{SiN}$ over a frequency rang of 7 kHz, making the probe tone to move out of resonance. However, the probe tone can still be amplified through constructive interference, since this frequency shift lies within $\gamma_{Al}$. When probing the Al drum, although the probe tone always remains within the bandwidth, clear constructive interference can be observed only when the resonance shift relative to the value calibrated at $V_{pc}$ lies within $\gamma_{SiN}$. We also noticed that the detected amplitudes of Al drum exhibiting small fluctuations for $V_{p}<$ 75 mV. It may be induced by fluctuations in the background noise carried by the pump tone. Fig. 2 (d) presents the calculation results, based on Eq.2, including the pump induced shift of $\Omega_{SiN}$, which are consistent with the experimental observations. These results demonstrate that pump amplitude modulation in the two-tone driving scheme provides a controllable way of inducing nonlinear behavior in the probe tone, in coupled resonators. In the following part, we exploit this nonlinear dynamics for reservoir computing.